OpenStudy (anonymous):

I give medals. c: (/^.^\) 8th grade math here!!

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Which shows this equation in standard form? 8 – 4(x – y) = 2x + 6 A. 4x – 6y = 2 B. 6x + 4y = 2 C. 6x – 4y = –2 D. 6x – 4y = 2

OpenStudy (anonymous):

what is it about math or medals?

OpenStudy (anonymous):

cant see half of it

OpenStudy (anonymous):

@CJDOG refresh the page

OpenStudy (anonymous):

Oh, sorry... il type it... Which shows this equation in standard form? 8 – 4(x – y) = 2x + 6

OpenStudy (anonymous):

ok thxs

OpenStudy (anonymous):

Ok, to solve the equation 8 - 4(x - y) = 2x + 6, you have to first do distributive property. Then you have 8 - 4x - 4y = 2x + 6. Then, you take away 2x from each side to get 8 - 6x - 4y = 6. Then, you take away 8 from each side to get -6x - 4y = -2.

OpenStudy (anonymous):

Also, this is all I need help with... Im in advanced, but im just stuck and my lesson doesnt provide enough information!!! gahhhhh So, it really isnt 8th grade... :/

OpenStudy (anonymous):

Never mind, I have the answer... sorry...

OpenStudy (anonymous):

Yeah, I'm in 7th grade and I'm doing 9th grade math.

OpenStudy (anonymous):

can anyone help with my question?

OpenStudy (anonymous):

I'm double advanced. XP

OpenStudy (anonymous):

Thank you for your help though!!!

OpenStudy (anonymous):

Welcome! :)

OpenStudy (anonymous):

Yay! Im doing 10th!! cx

OpenStudy (anonymous):

XP ur in 8th grade?

OpenStudy (anonymous):

can anyone come help me with a question?

OpenStudy (anonymous):

sure, what's the question? (or just post the link to it)

OpenStudy (anonymous):

Create a scenario that leads to an inequality of the form ax + b > c. You can use any of the inequality symbols (>, ≥, <, ≤) in your inequality. For this step, just write the word problem. Write the inequality and show all work to solve the inequality. Graph the solution to your inequality on a number line. Explain what your solution means in the context of the problem.

OpenStudy (anonymous):

It was b, if anyone else was wondering

OpenStudy (anonymous):

i just need a word problem

OpenStudy (anonymous):

Sorry, I can't answer that good. We're still learning about that.

OpenStudy (anonymous):

ok its 7th grade advanced

OpenStudy (anonymous):

*answer that question.

OpenStudy (anonymous):

cya

OpenStudy (anonymous):

cya!

OpenStudy (anonymous):

im a fan

OpenStudy (anonymous):

i saw dat